# Describing the Non-Universal Galaxy Merger Timescales in IllustrisTNG: Effects of Host Halo Mass, Baryons, and Sample Selection

### v1

@@ 5.2 The whole infall sample and selection effects
One remaining concern regarding our merger timescale results is selection effects, a common issue in studies of merger timescales using cosmological simulations. Since we only analyze satellites that have already merged with centrals, our sample is likely biased toward events with shorter merger timescales. If this sample is taken as representative of all infall events, our model may underestimate the true average merger timescales. With the model for arbitrary infall radii, we can now utilize the entire infall sample—whether the satellites have merged, dispersed, or survived—to study selection effects. This is possible because our model allows us to predict merger timescales for any two points in an event’s timeline and compare the predicted time intervals with those from simulations. As a result, satellites do not need to have already merged with the central to be included in the analysis.
@@
To construct the infall sample, we select all satellites that have crossed $r_{\rm vir}$ after $z=4$, using only the mass selection criterion as before: $M_{\rm{vir}}(z=0)>10^{10.7}h^{-1}M_{\odot}$ and $M_{*}^{\rm sat}(z_{\rm infall})>10^{8.3}h^{-1}M_{\odot}$ for TNG100-1, and $M_{\rm{vir}}(z=0)>10^{11.7}h^{-1}M_{\odot}$ and $M_{*}^{\rm sat}(z_{\rm infall})>10^{9.3}h^{-1}M_{\odot}$ for TNG300-1. In Figure 15, we present the distributions of the initial properties—$\log_{10}(M_{\rm sat}/M_{\rm host})$, $\epsilon$, and $r_{c}/r_{\rm host}$—at the time of first crossing $r_{\rm vir}$ for the infall sample. We also show the contributions from three sub-samples: satellites that have merged with centrals, those that have already been disrupted (defined as having $\log_{10}(j_{\rm merger}/j_{\rm infall})>-1$ when they disappeared), and those that still survive at $z=0$. If we use only the merger sample instead of the full infall sample as before, we tend to select satellites with higher mass ratios, more radial orbits , and possibly shorter merger timescale events for the same initial conditions, which may bias our model toward underestimating merger timescales.
@@
Therefore, we refine our model by incorporating the entire infall sample. For the merger sample, we maintain the previous approach, requiring only the prediction of the merger timescale at the infall time, $T_{\rm merger}(t_{\rm infall})$, where $t_{\rm infall}$ is defined as the time of first crossing $r_{\rm vir}$. For the disruption and surviving samples, we predict both $T_{\rm merger}(t_{\rm infall})$ and $T_{\rm merger}(t_{50})$, where $t_{50}$ is the last time the subhalo contained more than 50 particles. If the subhalo still has more than 50 particles at $z=0$, then $t_{50}$ corresponds to the present time. The requirement of at least 50 remaining particles is imposed to reduce numerical artifacts. With these predictions, we compute $\Delta T_{\rm merger}=T_{\rm merger}(t_{\rm infall})-T_{\rm merger}(t_{50})$ from our model and compare it with simulation results. For the merger sample, $\Delta T_{\rm merger}$ simply equals $T_{\rm merger}(t_{\rm infall})$.
@@
We first assess whether our previously derived arbitrary infall radii model, based on the merger sample, accurately predicts $\Delta T_{\rm merger}$ for the entire infall sample. Figure 16 compares model predictions of $\Delta T_{\rm merger}$ with simulation results. Overall, the model performs well across all three sub-samples, except when the mass ratio is very small ($M_{\rm sat}/M_{\rm host}<10^{-3}$). For the merger sample, $T_{\rm merger}$ from simulations is generally slightly shorter than the model predictions, with discrepancies reaching  0.2 dex for $M_{\rm sat}/M_{\rm host}<10^{-2}$. This aligns with expectations, as the merger sample tends to favor shorter merger timescales. Despite being calibrated on the merger sample, the model’s ability to incorporate information from arbitrary radii appears to mitigate selection effects, likely due to the collisional nature of the merger process.
@@
We refine our model by fitting $\Delta T_{\rm merger}$ from the entire infall sample, using the same framework as the arbitrary infall radii model. The best-fit parameters are listed in the fourth row of Table 1, and Figure 17 compares model predictions with simulation results. The new model shows improved accuracy, particularly for low mass ratios ($M_{\rm sat}/M_{\rm host}<10^{-3}$) and high $r_{c}/r$, with an average logarithmic scatter of $\sigma_{\log}=0.317$ for the entire infall sample. For $M_{\rm sat}/M_{\rm host}<10^{-2}$, $T_{\rm merger}$ from the merger sample remains 0.2 dex lower than the model prediction, highlighting the presence of selection effects. This also indicates that incorporating the entire infall sample into the fitting framework effectively reduces selection effects. However, this comes at the cost of increased scatter due to the need for predictions at two different times.
@@ 6 Conclusion and Discussion
  • Table 1 presents four best-fit models for merger timescales, each derived from different samples. We recommend using the fourth model, which is based on the entire infall sample and applicable to any infall radius. The other models may still be useful for specific cases.
@@
Our models for galaxy merger timescales, presented in Table 1, serve as valuable tools for semi-analytical and empirical models of galaxy formation. The evolution of subhalos (satellites) is a crucial component of hierarchical structure formation. Unlike halo formation and evolution—where physically motivated semi-analytical models such as the Press-Schechter framework (Press & Schechter, 1974; Bond et al., 1991; Lacey & Cole, 1993) provide theoretical guidance—subhalo evolution after halo mergers is highly non-linear and chaotic, making it challenging to model without numerical experiments. The merger timescale is a key parameter in determining subhalo abundance, making our models an essential input for semi-analytical and empirical approaches to galaxy formation. Moreover, even in high-resolution numerical simulations, halos are more resilient to numerical artifacts than subhalos. The abundance and distribution of halos can converge with as few as 10–20 particles (Jing, 2019). However, for subhalos, the high-density environment within host halos leads to rapid tidal disruption, requiring at least 50 particles for a robust simulation of subhalo evolution (van den Bosch & Ogiya, 2018; He et al., 2025), which can be computationally inefficient. In practice, the most bound particles of artificially disrupted subhalos are often tracked to approximate their unresolved evolution, but their survival time must be estimated carefully, as these particles may experience significantly weaker dynamical friction than fully resolved subhalos. If merger timescales are not considered when determining subhalo survival, their abundance may be overestimated. Thus, our galaxy merger timescale models play a crucial role in improving the accuracy of galaxy formation modeling.
Moreover, our merger timescale model is valuable for inferring merger rates from observed close pairs (Kitzbichler & White, 2008; Jiang et al., 2014), which is crucial for the ex situ formation of galaxies.
@@ Acknowledgments
We thank Jiaxin Han, Shaun Cole, and Carlos Frenk for their valuable discussions. We also appreciate Dylan Nelson and Vicente Rodriguez-Gomez for their assistance in accessing the merger trees of IllustrisTNG. K.X. is supported by the funding from the Center for Particle Cosmology at U Penn. Y.P.J. is supported by NSFC (12133006, 11890691), by National Key R&D Program of China (2023YFA1607800, 2023YFA1607801), by grant No. CMS-CSST-2021-A03, and by 111 project No. B20019. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.

### v4

@@ 5.2 The entire infall sample and selection effects
One remaining concern regarding our merger timescale results is selection effects, a common issue in studies of merger timescales using cosmological simulations. Since we only analyze satellites that have already merged with centrals, our sample is likely biased toward events with shorter merger timescales. If this sample is taken as representative of all infall events, our model may underestimate the true average merger timescales. With the model for arbitrary starting radii, we can now use the entire infall sample—whether the satellites have merged, been disrupted, or are still surviving—to study selection effects. This is possible because our model allows us to predict merger timescales for any two points in an event’s timeline and compare the predicted time intervals with those from simulations. As a result, satellites do not need to have already merged with the central to be included in the analysis.
@@
To construct the infall sample, we select all satellites that have crossed $r_{\rm vir}$ after $z=4$, using only the mass selection criterion as before: $M_{\rm{vir}}(z=0)>10^{10.7}h^{-1}M_{\odot}$ and $M_{*}^{\rm sat}(z_{\rm infall})>10^{8.3}h^{-1}M_{\odot}$ for TNG100-1, and $M_{\rm{vir}}(z=0)>10^{11.7}h^{-1}M_{\odot}$ and $M_{*}^{\rm sat}(z_{\rm infall})>10^{9.3}h^{-1}M_{\odot}$ for TNG300-1. In Figure 11, we present the distributions of the initial properties—$\log_{10}(M_{\rm sat}/M_{\rm host})$, $\epsilon$, and $r_{c}/r_{\rm host}$—at the time of first crossing $r_{\rm vir}$ for the infall sample. We also show the contributions from three sub-samples: satellites that have merged with the centrals ($\log_{10}(j_{\rm merger}/j_{\rm infall})<-1$), those that have been disrupted ($\log_{10}(j_{\rm merger}/j_{\rm infall})>-1$, regardless of whether the disruption is physical or artificial), and those that are still surviving at $z=0$. If we use only the merger sample instead of the full infall sample as before, we tend to select satellites with higher mass ratios, more radial orbits , and possibly shorter merger timescale events for the same initial conditions, which may bias our model toward underestimating merger timescales.
@@
Therefore, we refine our model by incorporating the entire infall sample. For the merger sample, we maintain the previous approach, requiring only the prediction of the merger timescale at the infall time, $T_{\rm merger}(t_{\rm infall})$, where $t_{\rm infall}$ is defined as the time of first crossing $r_{\rm vir}$. For the disruption and surviving samples, we predict both $T_{\rm merger}(t_{\rm infall})$ and $T_{\rm merger}(t_{50})$, where $t_{50}$ is the last time the subhalo contained more than 50 particles. If the subhalo still has more than 50 particles at $z=0$, then $t_{50}$ corresponds to the present time. The requirement of at least 50 remaining particles is imposed to reduce numerical artifacts. With these predictions, we compute $\Delta T_{\rm merger}=T_{\rm merger}(t_{\rm infall})-T_{\rm merger}(t_{50})$ from our model and compare it with simulation results. For the merger sample, $\Delta T_{\rm merger}$ simply equals $T_{\rm merger}(t_{\rm infall})$. We have also tested $t_{100}$ and found similar results, confirming the robustness of this choice.
@@
We first assess whether our previously derived arbitrary starting radii model, based on the merger sample, accurately predicts $\Delta T_{\rm merger}$ for the entire infall sample. Figure 12 compares model predictions of $\Delta T_{\rm merger}$ with simulation results. Overall, the model performs well across all three sub-samples, except when the mass ratio is very small ($M_{\rm sat}/M_{\rm host}<10^{-3}$). For the merger sample, $T_{\rm merger}$ from simulations is generally slightly shorter than the model predictions, with discrepancies reaching  0.2 dex for $M_{\rm sat}/M_{\rm host}<10^{-2}$. This aligns with expectations, as the merger sample tends to favor shorter merger timescales. Despite being calibrated on the merger sample, the model’s ability to incorporate information from arbitrary radii appears to mitigate selection effects, likely due to the collisional nature of the merger process.
@@
We refine our model by fitting $\Delta T_{\rm merger}$ from the entire infall sample, using the same framework as the arbitrary starting radii model. The best-fit parameters are listed in the fourth row of Table 1, and Figure 13 compares model predictions with simulation results. The new model shows improved accuracy, particularly for low mass ratios ($M_{\rm sat}/M_{\rm host}<10^{-3}$) and high $r_{c}/r$, with an average logarithmic scatter of $\sigma_{\log}=0.317$ for the entire infall sample. For $M_{\rm sat}/M_{\rm host}<10^{-2}$, $T_{\rm merger}$ from the merger sample remains 0.2 dex lower than the model prediction, highlighting the presence of selection effects. This also indicates that incorporating the entire infall sample into the fitting framework effectively reduces selection effects. However, this comes at the cost of increased scatter due to the need for predictions at two different times.
@@ 6 Conclusion and Discussion
  • Table 1 presents four best-fit models for merger timescales, each derived from different samples. We recommend using the fourth model, which is based on the entire infall sample and applicable to any starting radius. The other models may still be useful for specific cases.
@@
Our models for galaxy merger timescales, presented in Table 1, serve as valuable tools for semi-analytical and empirical models of galaxy formation. The evolution of subhalos (satellites) is a crucial component of hierarchical structure formation. Unlike halo formation and evolution—where physically motivated semi-analytical models such as the Press-Schechter framework (Press & Schechter, 1974; Bond et al., 1991; Lacey & Cole, 1993) provide theoretical guidance—subhalo evolution after halo mergers is highly non-linear and chaotic, making it challenging to model without numerical experiments (Taylor & Babul, 2001; Kazantzidis et al., 2004; Han et al., 2016; Drakos et al., 2022; Stücker et al., 2023). The merger timescale is a key parameter in determining subhalo abundance, making our models an essential input for semi-analytical and empirical approaches to galaxy formation. Moreover, even in high-resolution numerical simulations, halos are more resilient to numerical artifacts than subhalos. The abundance and distribution of halos can converge with as few as 10–20 particles (Jing, 2019). However, for subhalos, the high-density environment within host halos leads to rapid tidal disruption, requiring at least 50 particles for a robust simulation of subhalo evolution (van den Bosch & Ogiya, 2018; He et al., 2025), which can be computationally inefficient. In practice, the most bound particles of artificially disrupted subhalos are often tracked to approximate their unresolved evolution, but their survival time must be estimated carefully, as these particles may experience significantly weaker dynamical friction than fully resolved subhalos. If merger timescales are not considered when determining subhalo survival, their abundance may be overestimated. Thus, our galaxy merger timescale models play a crucial role in improving the accuracy of galaxy formation modeling.
Our merger timescale model is also useful for inferring galaxy merger rates from observed close pairs (Kitzbichler & White, 2008; Jiang et al., 2014), which play a key role in the ex situ growth of galaxies. Estimating galaxy merger rates typically involves identifying galaxy pairs and determining merger timescales as a function of separation distance, where our model with arbitrary starting radii can provide valuable input (Ferreira et al., 2020; Whitney et al., 2021). Moreover, the galaxy merger timescales defined in our study effectively represent the delay between halo mergers and the corresponding galaxy mergers. As such, they can be useful for inferring the halo merger rate, which is not directly observable but is closely linked to cosmological parameters (Amoura et al., 2024).
@@ Acknowledgments
We thank the anonymous referee for the detailed and valuable comments that helped improve the manuscript. We thank Jiaxin Han, Shaun Cole, and Carlos Frenk for their valuable discussions. We also appreciate Dylan Nelson and Vicente Rodriguez-Gomez for their assistance in accessing the merger trees of IllustrisTNG. K.X. is supported by the funding from the Center for Particle Cosmology at U Penn. Y.P.J. is supported by NSFC (12133006, 11890691), by National Key R&D Program of China (2023YFA1607800, 2023YFA1607801), by grant No. CMS-CSST-2021-A03, and by 111 project No. B20019. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
[…]
Appendix A Validation of fits
In this section, we validate the merger timescale fits across different $M_{\rm host}$ bins. Results for TNG100-1, TNG300-1, TNG100-1-Dark, and TNG300-1-Dark are shown in Figures 14, 15, 16, and 17, respectively.